### v1

@@ I Introduction
One method to solve this problem is the neutron DFI [19], which uses neutrons instead of laser light. Since neutrons can, in principle, have arbitrary speeds, the DFI’s effective frequency can be reduced using neutrons with speeds slower than light. Another advantage of the neutron DFI is the simplification of the interferometer configuration. Previous studies have demonstrated that the number of optical components and the required neutron incidence angles can be reduced or optimized using multiple neutron speeds [20, 21], with further refinements explored in [22].
@@ II Review of Neutron DFI using four-speed unidirectional neutrons
Let $T_{\mathrm{i}}$ represent the propagation time of a neutron along one side of the interferometer square. When neutrons at different speeds reach mirrors simultaneously, the associated phase shifts at $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ are given by $\phi_{\text{BSU},\mathrm{i}}(t-T_{\mathrm{i}})$ and $\phi_{\text{BSL},\mathrm{i}}(t+T_{\mathrm{i}})$, respectively. For GWs, we assume the neutron phase shift occurs along one leg of the interferometer path, around the midpoint between a mirror and a beam splitter (BS). This simplification yields GW-induced phase shifts of $\phi_{\text{GWU},\mathrm{i}}(t-T_{i}/2)$ and $\phi_{\text{GWL},\mathrm{i}}(t+T_{\mathrm{i}}/2$). Accordingly, the total phase shift for each neutron speed is: $\phi_{\mathrm{i}}(t)=\phi_{\mathrm{BS_{U},\mathrm{i}}}(t-T_{\mathrm{i}})+\phi_{\mathrm{GW_{U},\mathrm{i}}}(t-T_{\mathrm{i}}/2)+\phi_{\mathrm{M_{1},\mathrm{i}}}(t)+\phi_{\mathrm{M_{2},i}}(t)+\phi_{\mathrm{GW_{L},i}}(t+T_{\mathrm{i}}/2)+\phi_{\mathrm{BS_{L},i}}(t+T_{\mathrm{i}})$, as shown in Table 1. Here, we assume that all terms carry a positive sign for simplicity.
@@
[…]
@@
This condition makes it possible to cancel the displacement noises of $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ with the DFI signal combination, which is written as
@@
The key to canceling the $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ noise components in Fig. 2 — after the mirror displacement noise has been removed — lies in the fact that the tangerine and purple arrows are parallel. The $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ noise can be eliminated by subtracting these two vectors, weighted by coefficients $c_{14}(\Omega)$ and $c_{23}(\Omega)$ that account for their amplitude differences. In contrast to BS noises, the residual GW signals are preserved in Eq. (4) because the length ratio of the arrows in the GW signals in Fig. 3 is different from that in BS noises.
@@ III Comparison between ideal and actual experiment
In this ideal configuration, both simulated displacement noise and GW signals are introduced, enabling validation of the DFI principle: appropriate combinations of interferometric outputs can cancel displacement noise while preserving GW signals. The neutron beams are assumed to strike mirrors simultaneously, and the interferometer’s arm lengths and modulation frequencies are selected to ensure sufficient phase differences, typically on the order of $\pi$ between neutrons of different speeds. As established in Section II, this phase difference corresponds to angular separation in the phasor diagram, which is essential for differentiating noise from GW signals. Therefore, the experimental setup must be designed with appropriate modulation depth for the simulated signal and noise and a suitable path length of the interferometer.
@@
However, our experimental setup deviates from this ideal due to several practical constraints. As shown in Fig. III, the experiment was carried out at the Materials and Life Science Experimental Facility (MLF) at J-PARC [26], using equipment based on an earlier neutron interferometer demonstration [28] that successfully employed beam-splitting etalons (BSEs).
[…]
figureActual setup for the neutron DFI used at J-PARC (left panel), and a photograph of the setup with the inserted sample displayed (right panel). In the left panel, neutron mirrors and BSs correspond to magenta lines and the paths of interfering neutrons are illustrated by the solid (Path 1) and dashed (Path 2) black arrows. In the photograph of the right panel, two beam-splitting etalons (BSEs) are highlighted by yellow squares, and the neutron beam enters there from the top side of the picture.
[…]
Another significant limitation lies in the available path length. Achieving sufficient phase modulation requires a long path relative to the modulation frequency. However, the practical path length is limited to approximately 0.2 meters, constrained by alignment tolerances of the BSEs. This short path length limits the achievable phase difference between neutrons of different speeds to below $\pi$.
@@ IV.1 Experimental Apparatus
[…]
figureActual setup for the proof-of-principle experiment (left panel) and photograph of an inserted sample (right panel). In the picture in the right panel, a pair of Al phase plates of modulator and compensator is highlighted by the yellow square and the actuators (piezoelectric transducers) are indicated by the green arrows.
@@ IV.3 Demonstration Setup and Analysis for the DFI
Four neutron speeds were selected at BL05: $v_{1}=467.3$ m/s, $v_{2}=435.0$ m/s, $v_{3}=381.7$ m/s, $v_{4}=360.0$ m/s, all of which satisfy the DFI velocity condition defined in Eq. 3. In practice, these neutron speeds are identified using time-of-flight (TOF) measurements, which typically carry an uncertainty of $\pm$50 $\mu$s — corresponding to a velocity error of approximately $\pm$0.5 $m$/s. This effect should be considered in the analysis of experimental results. For the simulated data, we consider five types of signals: $\mathrm{BS_{U}}$, $\mathrm{GW_{U}}$, $\mathrm{M}$ (mirror noise), $\mathrm{GW_{L}}$, and $\mathrm{BS_{L}}$. Among these, the mirror noise signal is assigned a frequency of $f_{\mathrm{M}}=200$ Hz, which serves as the time reference for canceling the displacement noise originating from the BSs.
@@
where $I_{\mathrm{path1}}$ and $I_{\mathrm{path2}}$ are the neutron counts for the beams of path 1 and path 2 (shown in Fig. III), respectively, and $I_{\mathrm{path1},\mathrm{SinglePass}}$ and $I_{\mathrm{path2},\mathrm{SinglePass}}$ are the neutron counts for each path when the other path is blocked by a Cd slit [28]. Thus, $I(\lambda)$ ranges at most from -1 to +1, depending on neutron’s wavelength (and thus the TOF). It is important to note that the measured interference fringes, a function of neutron wavelength, depend on the optical path length offset within the interferometer. The modulation signals discussed later in this section are derived from these measured interference fringes.
@@
The modulation signals as a function of the TOF of neutrons and the initial phase of the simulated signal, as well as those as a function of the initial phase at the TOF of each neutron for the five kinds of modulation signals, are shown in Figs. 9-12.
@@
Next, the compensation methods described in Fig. 8 are applied to the GW signals and the BS noises. Specifically, each curve in the right panel of Figs. 9-12 is shifted by an amount specified by the arrows in Fig. 8. The modulation signals as a function of the initial phase after compensation are shown in Figs. 13-14.
@@
After applying the same phase compensation, the phasor diagrams for $\mathrm{BS_{U}}$, $\mathrm{GW_{U}}$, $\mathrm{GW_{L}}$, and $\mathrm{BS_{L}}$, corresponding to Figs. 13-14 are shown in Figs. 16-17. These figures also display the residual signals after all mirror noises are canceled as described in Eqs. (1)-(2). It should be noted that the length ratio between the tangerine and purple arrows (which are parallel in each panel) for $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ is equal, and the same holds for $\mathrm{GW_{U}}$ and $\mathrm{GW_{L}}$. However, the length ratio for $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ differs from that for $\mathrm{GW_{U}}$ and $\mathrm{GW_{L}}$.
@@ V Experimental results
$+P_{\mathrm{AC}}\sin\bigl(2\pi f_{\mathrm{AC}}t+\phi_{\mathrm{initial}}\bigr{missing})\Bigr{)}\lambda\Bigr{\}}+B,$ (16)
where $A$ is the visibility of the interference fringe, $B$ is the offset of the interference fringe, $P_{L}$ is the geometric optical length difference, $P_{\mathrm{IFO}}$ is the path difference from the $\mathrm{SiO_{2}}$ BSEs of the interferometer, $P_{\mathrm{sample}}$ is the path difference from a sample without modulation, $P_{\mathrm{offset}}$ is the offset of the path difference from a modulated sample, $P_{\mathrm{AC}}$ is the path difference from the modulation amplitude, $f_{\mathrm{AC}}$ is the modulation frequency, and $\phi_{\mathrm{initial}}$ is the initial phase of the modulation signal, which is used for the phase scan. The details of the fitting process are denoted in Appendix Appendix: Fitting process for experimental results. The fitting results, including 100 Hz, 150 Hz, 250 Hz, and 300 Hz measurements, are shown in Table 4.
@@
Frequency $P_{\mathrm{AC}}$ $[\mathrm{nm}^{-1}]$ $f_{\mathrm{AC}}$ $[\mathrm{Hz}]$
@@
Using these parameters, the modulation signals can be obtained through a similar process to that in Section IV.3. For the mirror noise modulation, the simulation data in Figs. 7-8 are used as a time reference. The interference fringes as a function of the TOF of neutrons and the initial phase of the simulated signal, as well as the interference fringes as a function of the TOF of neutrons, for $\mathrm{BS_{U}}$, $\mathrm{GW_{U}}$, $\mathrm{GW_{L}}$, and $\mathrm{BS_{L}}$ obtained from measurements are shown in Figs. 21-24. Notably, the measured modulation signals agree with the simulated modulation signals shown in Figs. 9-12 within the margin of error, except for a slight difference in the overall magnitude.
@@
In Figs. 21-24, the curves representing modulation signals are obtained by fitting the data with a sine function. Subsequently, the same compensation method as in Fig. 8 is applied to the GW signals and the BS noises, as shown in Figs. 25-26. The measured modulation signals after compensation also agree with the simulated modulation signals shown in Figs. 13-14 within the margin of error, except for a slight difference in the overall magnitude.
@@
Figs. V-V show the measured phasor diagrams corresponding to the $\mathrm{BS_{U}}$, $\mathrm{BS_{U}}$, $\mathrm{GW_{U}}$, and $\mathrm{GW_{L}}$ signals. These diagrams also illustrate the residual signals after cancellation of all mirror noises. The measured phasor diagram agrees with the simulated phasor diagram shown in Figs. 16-17, including the relationships between the arrows, within the margin of error, except for a slight difference in the overall magnitude.
[…]
figureMeasured phasor diagram of measured $\mathrm{BS_{U}}$ (upper panel) and $\mathrm{BS_{L}}$ (lower panel) displacement noises, respectively. The red, blue, green, and magenta arrows show the neutron speeds of $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, respectively. The tangerine and purple arrows indicate the signals defined in Eqs. (1) and (2). The shaded regions represent their errors in amplitude and phase, which are colored according to their associated arrows. The dashed circle indicates the amplitude of the simulation result.
[…]
figureMeasured phasor diagram of measured $\mathrm{GW_{U}}$ (upper panel) and $\mathrm{GW_{L}}$ (lower panel) signals. All parameters and definitions in this figure are the same as those in Fig. V.
@@
Fig. 27 shows the experimental result demonstrating the dependence of the GW signals and BS noises in the DFI combination signal on the coefficient ratio is shown, along with the error regions. This result shows that BS noises are canceled at the correct coefficient ratio within the margin of error, while preserving the GW signal.
@@
Fig. 28 shows the dependence of the BS noise and GW signal in the DFI combination signal on the coefficient ratio is shown, along with the error regions. The figure shows that the coefficient ratio that cancels out the BS noise matches the predicted value ($c_{1}4/c_{2}3\sim-4.31$) from the simulation within the margin of error. Importantly, despite the errors, the GW signal is unequivocally preserved at this correct coefficient ratio.

### v3

@@ I Introduction
One method to solve this problem is the neutron DFI [19], which uses neutrons instead of laser light. Since neutron speed can be chosen within experimental constraints, the DFI’s effective frequency can be reduced using neutrons with speeds slower than light. Another advantage of the neutron DFI is the simplification of the interferometer configuration. Previous studies have demonstrated that the number of optical components and the required neutron incidence angles can be reduced or optimized using multiple neutron speeds [20, 21], with further refinements explored in [22].
@@ II Review of Neutron DFI using four-speed unidirectional neutrons
Let $T_{\mathrm{i}}$ represent the propagation time of a neutron along one side of the interferometer square. When neutrons at different speeds reach mirrors simultaneously, the associated phase shifts at $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ are given by $\phi_{\text{BSU},\mathrm{i}}(t-T_{\mathrm{i}})$ and $\phi_{\text{BSL},\mathrm{i}}(t+T_{\mathrm{i}})$, respectively. For GWs, we assume the neutron phase shift occurs along one leg of the interferometer path, around the midpoint between a mirror and a beam splitter (BS). This simplification yields GW-induced phase shifts of $\phi_{\text{GWU},\mathrm{i}}(t-T_{i}/2)$ and $\phi_{\text{GWL},\mathrm{i}}(t+T_{\mathrm{i}}/2$). Accordingly, the total phase shift for each neutron speed is: $\phi_{\mathrm{i}}(t)=\phi_{\mathrm{BS_{U},\mathrm{i}}}(t-T_{\mathrm{i}})+\phi_{\mathrm{GW_{U},\mathrm{i}}}(t-T_{\mathrm{i}}/2)+\phi_{\mathrm{M_{1},\mathrm{i}}}(t)+\phi_{\mathrm{M_{2},i}}(t)+\phi_{\mathrm{GW_{L},i}}(t+T_{\mathrm{i}}/2)+\phi_{\mathrm{BS_{L},i}}(t+T_{\mathrm{i}})$, as shown in Table 1.
@@
Here, we assume that all terms carry a positive sign for simplicity. This is justified because in both paths the beams experience a reflection on the same side of the beam splitter, resulting in a common phase shift that cancels out.
@@
Now, we consider the interference signal in the frequency domain instead of the time domain to explain the cancellation of the displacement noises from $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$. The Fourier spectrum of the interference signal is defined as $S_{\mathrm{x}}(\Omega)=\int_{-\infty}^{\infty}s_{\mathrm{x}}(t)e^{-j\Omega t}dt$, where $\Omega$ is the angular frequency, $j$ is an imaginary unit and x is arbitrary, representing i or a pair of i and l. Accordingly, the time-domain interference signal can be expressed as $s_{\mathrm{x}}(t)=\int_{-\infty}^{\infty}S_{\mathrm{x}}(\Omega)e^{j\Omega t}\,d\Omega$. Here, we impose a condition on the neutron speeds, which is given by the following equation (see [21] for more details; an intuitive explanation is provided later in this section).
@@
This is based on selecting a speed such that the average of phase between $v_{1}$ and $v_{4}$ is the same as that between $v_{2}$ and $v_{3}$. This condition makes it possible to cancel the displacement noises of $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ with the DFI signal combination, which is written as
@@
The key to canceling the $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ noise components in Fig. 2 — after the mirror displacement noise has been removed — lies in the fact that the tangerine and purple arrows are parallel. The $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ noise can be eliminated by subtracting these two vectors, weighted by coefficients $c_{14}(\Omega)$ and $c_{23}(\Omega)$ that account for their amplitude differences. In contrast to BS noises, the residual GW signals are preserved in Eq. (4) because the length ratio of the arrows in the GW signals in Fig. 3 is different from that in BS noises. In Fig. 2, the tangerine arrow is more than twice as long as the purple one and points in the opposite direction. In contrast, in Fig. 3, the tangerine arrow is slightly shorter than the purple one and points in the same direction. This difference illustrates that the GW signals are not canceled by the combination that eliminates the BS noises.
@@ III Comparison between ideal and actual experiment
In this ideal configuration, both simulated displacement noise and GW signals are introduced, enabling validation of the DFI principle: appropriate combinations of interferometric outputs can cancel displacement noise while preserving GW signals. The neutron beams are assumed to strike mirrors simultaneously, which allows the mirror noise cancellation in Eqs. (1)-(2), as discussed in Section II, and the interferometer’s arm lengths and modulation frequencies are selected to ensure sufficient phase differences, typically on the order of $\pi$ between neutrons of different speeds. As established in Section II, this phase difference corresponds to angular separation in the phasor diagram, which is essential for differentiating noise from GW signals. Therefore, the experimental setup must be designed with appropriate modulation depth for the simulated signal and noise and a suitable path length of the interferometer.
@@
However, our experimental setup deviates from this ideal due to several practical constraints. As shown in Fig. 5, the experiment was carried out at the Materials and Life Science Experimental Facility (MLF) at J-PARC [26], using equipment based on an earlier neutron interferometer demonstration [28] that successfully employed beam-splitting etalons (BSEs).
Figure 5: Actual setup for the neutron DFI used at J-PARC (left panel), and a photograph of the setup with the inserted sample displayed (right panel). In the left panel, neutron mirrors and BSs correspond to magenta lines and the paths of interfering neutrons are illustrated by the solid (Path 1) and dashed (Path 2) black arrows. In the photograph of the right panel, two beam-splitting etalons (BSEs) are highlighted by yellow squares, and the neutron beam enters there from the top side of the picture.
The interferometer adopts a parallelogram geometry with an acute angle of approximately $1\tcdegree$, instead of a square. Instead of separate mirrors and BSs with independent actuators, two BSEs are used [23, 24]. These BSEs consist of two super-polished $\mathrm{SiO_{2}}$ plates facing each other with fixed spacers. Consequently, the mirrors and BSs are physically integrated and cannot be actuated independently, unlike in the ideal setup.
Another significant limitation lies in the available path length. Achieving sufficient phase modulation requires a long path relative to the modulation frequency; for example, a path length of approximately 1.25 m is required to obtain half-wave phase modulation at a modulation frequency of 200 Hz for neutrons traveling at 500 m/s. However, the practical path length is limited to approximately 0.2 meters, constrained by alignment tolerances of the BSEs. This short path length limits the achievable phase difference between neutrons of different speeds to below $\pi$.
@@ IV.1 Experimental Apparatus
Fig. 6 shows the experimental layout at J-PARC. A pair of aluminum (Al) phase plates—used to induce neutron phase shifts—is inserted at a $45\tcdegree$ angle into both neutron paths. The pair consists of a modulator and a compensator. The compensator offsets the beam path displacement introduced by the modulator. Each Al plate features a zigzag shape to reduce spatial requirements while maintaining the $45\tcdegree$ insertion angle. The thickness of the plates along the beam path is 6 mm.
Figure 6: Actual setup for the proof-of-principle experiment (left panel) and photograph of an inserted sample (right panel). In the picture in the right panel, a pair of Al phase plates of modulator and compensator is highlighted by the yellow square and the actuators (piezoelectric transducers) are indicated by the green arrows.
@@ IV.3 Demonstration Setup and Analysis for the DFI
[…]
@@
where $I_{\mathrm{path1}}$ and $I_{\mathrm{path2}}$ are the neutron counts for the beams of path 1 and path 2 (shown in Fig. 5), respectively, and $I_{\mathrm{path1},\mathrm{SinglePass}}$ and $I_{\mathrm{path2},\mathrm{SinglePass}}$ are the neutron counts for each path when the other path is blocked by a Cd slit [28]. Thus, $I(\lambda)$ ranges at most from -1 to +1, depending on neutron’s wavelength (and thus the TOF). It is important to note that the measured interference fringes, a function of neutron wavelength, depend on the optical path length offset within the interferometer. The modulation signals discussed later in this section are derived from these measured interference fringes.
@@
The modulation signals as a function of the TOF of neutrons and the initial phase of the simulated signal, as well as those as a function of the initial phase at the TOF of each neutron for the five kinds of modulation signals, are shown in Figs. 11-14.
@@
Next, the compensation methods described in Fig. 10 are applied to the GW signals and the BS noises. Specifically, each curve in the right panel of Figs. 11-14 is shifted by an amount specified by the arrows in Fig. 10. The modulation signals as a function of the initial phase after compensation are shown in Figs. 15-16.
@@
After applying the same phase compensation, the phasor diagrams for $\mathrm{BS_{U}}$, $\mathrm{GW_{U}}$, $\mathrm{GW_{L}}$, and $\mathrm{BS_{L}}$, corresponding to Figs. 15-16 are shown in Figs. 18-19. These figures also display the residual signals after all mirror noises are canceled as described in Eqs. (1)-(2). It should be noted that the length ratio between the tangerine and purple arrows (which are parallel in each panel) for $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ is equal, and the same holds for $\mathrm{GW_{U}}$ and $\mathrm{GW_{L}}$. However, the length ratio for $\mathrm{BS_{U}}$ and $\mathrm{BS_{L}}$ differs from that for $\mathrm{GW_{U}}$ and $\mathrm{GW_{L}}$.
@@ V Experimental results
$+P_{\mathrm{AC}}\sin\bigl(2\pi f_{\mathrm{AC}}t+\phi_{\mathrm{initial}}\bigr.)\Bigr)\lambda\Bigr\}+B,$ (16)
where $A$ is the visibility of the interference fringe, $B$ is the offset of the interference fringe, $P_{L}$ is the geometric optical length difference, $P_{\mathrm{IFO}}$ is the path difference from the $\mathrm{SiO_{2}}$ BSEs of the interferometer, $P_{\mathrm{sample}}$ is the path difference from a sample without modulation, $P_{\mathrm{offset}}$ is the offset of the path difference from a modulated sample, $P_{\mathrm{AC}}$ is the path difference from the modulation amplitude, $f_{\mathrm{AC}}$ is the modulation frequency, and $\phi_{\mathrm{initial}}$ is the initial phase of the modulation signal, which is used for the phase scan. The details of the fitting process are denoted in Appendix A. The fitting results, including 100 Hz, 150 Hz, 250 Hz, and 300 Hz measurements, are shown in Table 4.
@@
Frequency $P_{\mathrm{AC}}$ $\mathrm{nm}^{-1}]$ $f_{\mathrm{AC}}$ $\mathrm{Hz}]$
@@
Using these parameters, the modulation signals can be obtained through a similar process to that in Section IV.3. For the mirror noise modulation, the simulation data in Figs. 9-10 are used as a time reference. The interference fringes as a function of the TOF of neutrons and the initial phase of the simulated signal, as well as the interference fringes as a function of the TOF of neutrons, for $\mathrm{BS_{U}}$, $\mathrm{GW_{U}}$, $\mathrm{GW_{L}}$, and $\mathrm{BS_{L}}$ obtained from measurements are shown in Figs. 23-26. Notably, the measured modulation signals agree with the simulated modulation signals shown in Figs. 11-14 within the margin of error, except for a slight difference in the overall magnitude.
@@
In Figs. 23-26, the curves representing modulation signals are obtained by fitting the data with a sine function. Subsequently, the same compensation method as in Fig. 10 is applied to the GW signals and the BS noises, as shown in Figs. 27-28. The measured modulation signals after compensation also agree with the simulated modulation signals shown in Figs. 15-16 within the margin of error, except for a slight difference in the overall magnitude.
@@
Figs. V-V show the measured phasor diagrams corresponding to the $\mathrm{BS_{U}}$, $\mathrm{BS_{L}}$, $\mathrm{GW_{U}}$, and $\mathrm{GW_{L}}$ signals. These diagrams also illustrate the residual signals after cancellation of all mirror noises. The measured phasor diagram agrees with the simulated phasor diagram shown in Figs. 18-19, including the relationships between the arrows, within the margin of error, except for a slight difference in the overall magnitude.
Figure 29: Measured phasor diagram of measured $\mathrm{BS_{U}}$ (upper panel) and $\mathrm{BS_{L}}$ (lower panel) displacement noises, respectively. The red, blue, green, and magenta arrows show the neutron speeds of $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, respectively. The tangerine and purple arrows indicate the signals defined in Eqs. (1) and (2). The shaded regions represent their errors in amplitude and phase, which are colored according to their associated arrows. The dashed circle indicates the amplitude of the simulation result.
Figure 30: Measured phasor diagram of measured $\mathrm{GW_{U}}$ (upper panel) and $\mathrm{GW_{L}}$ (lower panel) signals. All parameters and definitions in this figure are the same as those in Fig. V.
@@
[…]
@@
Fig. 32 shows the dependence of the BS noise and GW signal in the DFI combination signal on the coefficient ratio is shown, along with the error regions. The definition of shaded error bands in Fig. 32 is the same as that in Fig. 31. The figure shows that the coefficient ratio that cancels out the BS noise matches the predicted value ($c_{1}4/c_{2}3\sim-4.31$) from the simulation within the margin of error. Importantly, despite the errors, the GW signal is unequivocally preserved at this correct coefficient ratio.